Six men can build a 150 m wall in 10 hrs. How many men are needed to build a 200 m wall in 4 hrs?
Answer by richwmiller(17219)

6*10=60 man hours to build 150 m
60/150=2/5 man hours per m
200*2/5=80 man hours
80/4 =20 men